# Efficient and Generalizable Certified Unlearning: A Hessian-free Recollection Approach

### v3

@@ D.2 Impact of Epochs
Figure 4: Impact of Epoch. (a)(b) illustrate the in the distance metric for LR and CNN by varying epoch from 10 and 100 with 20% data to be forgotten. As shown in the Figure (a)(b), the error accumulates with the increase in epochs. However, with an appropriate choice of step size, such growth is entirely acceptable (Seed 42).
Impact of epochs: In Figure 4, we can observe that the distance slowly increases with the growth of epochs. This is attributed to the fact that each iteration involves an approximation, and multiple rounds of iterations lead to the accumulation on approximation error. However, as Figure 4 shows that, the accumulation of this error is small and does not lead to a significant performance degradation.
D.3 Impact of Decay Rates of Stepsize
This experiments aim to study the impacts of stepsizes. We consider two scenarios: one with stepsize decay ($q=0.995$) and the other without stepsize decay ($q=1$).
Impact of decay rates of stepsize: From Figure 5, we can observe that the Euclidean distance for the stepsize decay strategy is generally smaller compared to that without the stepsize decay. This aligns with our earlier analysis of step size in Subsection D.1, i.e., a smaller step size leads to a smaller approximation error and choosing an appropriate stepsize decay rate ensures a continuous reduction in step size, preventing the unbounded growth of the approximation error.
@@ D.3 Impact of Decay Rates of Stepsize
Figure 5: Impacts of the stepsize decay rate on the Euclidean distances for (a) the logistic regression and (b) the CNN by varying epochs from 1 to 200. The green line represents learning with a 0.995 stepsize decay rate, while the blue line represents without stepsize decay. It can be observed that stepsize decay effectively reduces the approximation error. (Seed 42)
Additionally, as in Figure 5 (b), we find that the distance of CNN tends to stabilize even without stepsize decay. This result can be explained by the conclusions in [53, 54]. Specifically, [53, 54] analyze the spectral distribution of the Hessian eigenvalues during CNN training. It reveals that, initially, the eigenvalues of Hessian are symmetrically distributed around 0. As training progresses, they converge towards 0, indicating Hessian degeneracy. Towards the end of training, only a small number of eigenvalues become negative and a large portion of the Hessian eigenvalues approach 0 when the training closes to convergence. Let’s recall our previous analysis of stepsize in Subsection D.1, where we demonstrates that the approximation error is primarily determined by $\rho$ of $\mathbf{I}-\frac{\eta_{e,b}}{|\mathcal{B}_{e,b}|}\mathbf{H}_{e,b}$, and $\rho$ is determined by $\eta_{e,b}$ and $\lambda_{1}$. As the number of iterations increases, the eigenvalues of the loss Hessian $\mathbf{H}_{e,b}$ tends to 0. This implies a decrease in $\lambda_{1}$, leading to a slower increase in error. Therefore, in Figure 5 (b), the distance gradually stabilizes even without a decay strategy due to the decrease in $\lambda_{1}$ during training, and it still increases due to the existence of some larger outlier eigenvalues.
D.4 Impact of Dataset Sizes
Finally, we assess the impacts of different dataset sizes. To maintain the total number of iterations, we adapt the number of epochs based on the dataset size. We keep other hyperparameters constant.
Table 10: Impact of Dataset Sizes on Convex Setting. Datasize $n=1,000$ represent hyperparameter choices of previous experiments for LR in our paper. (Seed 42)
Datasize Distance
1% 5% 20% 30%
500 0.043 0.087 0.212 0.300
1,000 0.030 0.060 0.162 0.230
5,000 0.014 0.036 0.104 0.1521
10,000 0.009 0.028 0.089 0.1334
25,000 0.007 0.024 0.081 0.1195
50,000 0.006 0.022 0.080 0.119
Table 11: Impact of Dataset Sizes on Non-Convex Setting. Datasize $n=1,000$ represent hyperparameter choices of previous experiments for CNN in our paper. (Seed 42)
Datasize Distance
1% 5% 20% 30%
500 0.108 0.237 0.592 0.837
1,000 0.167 0.380 0.584 0.825
2,000 0.092 0.230 0.483 0.715
5,000 0.081 0.177 0.447 0.677
10,000 0.132 0.191 0.463 0.654
20,000 0.165 0.190 0.460 0.617
Impact of dataset sizes: As shown in Table 11 and Table 11, we observe that larger dataset size generally tends to decrease the approximation error, making the unlearned model more closely resemble the retrained model. A reasonable explanation is that the enlargement of the dataset implies less frequent sampling of a specific sample for updating the model, consequently reducing the impact of data on the model updates in (3). This, in turn, leads to a decrease in the approximation error.
Appendix E Discussion
E.1 Deletion Capacity
We analyze the deletion capacity of our method based on the analysis of Sekhari et al. [2], which formalizes how many samples can be deleted from the model parameterized by the original empirical risk minimizer while maintaining reasonable guarantees on the test loss. We establish the relationship between generalization performance and the amount of deleted samples through deletion capacity, and our analysis indicates that our proposed method outperforms previous works in both generalization and deletion capacity. Below, we first formalize the deletion capacity.
Definition 5 (Definition of Deletion Capacity [2]).
Let $\varepsilon,\delta\geq 0$. Let $S$ be a dataset of size $n$ drawn i.i.d. from $\mathcal{D}$, and let $\ell(w;z)$ be a loss function. For a pair of learning and unlearning algorithms ${\Omega},\bar{\Omega}$ that are $(\varepsilon,\delta)$-unlearning, the deletion capacity $m_{\varepsilon,\delta}^{\Omega,\bar{\Omega}}(d,n)$ is defined as the maximum number of samples $U$ to be forgotten, while still ensuring an excess population risk is $\gamma$. Specifically,
$m_{\varepsilon,\delta,\gamma}^{\Omega,\bar{\Omega}}(d,n)$ $:=\max\left\{m\mid\mathbb{E}\left[\max_{U\subseteq\mathcal{S}:|U|\leq m}F(\bar{\Omega}(\Omega(\mathcal{S});T(\mathcal{S}))-F(\mathbf{w}^{*})\right]\leq\gamma\right\},$ (63)
Similar to Sekhari et al. [2], we provide upper bound and lower bound of deletion capacity for our algorithm. We first give the upper bound of deletion capacity.
Theorem 8 (Upper Bound of Deletion Capacity [2]).
$m_{\epsilon,\delta,\gamma}^{A,M}(d,n)\leq cn$ (64)
A comprehensive proof of Theorem 8 is available in Sekhari et al. [2]. Furthermore, based on Theorem 5, we present the lower bound of deletion capacity below:
Theorem 9 (Lower Bound of Deletion Capacity).
$m_{\varepsilon,\delta,\gamma}^{\Omega,\bar{\Omega}}(d,n)\geq n-\frac{4L^{2}}{\gamma\lambda}$ (65)
Proof of Theorem 9.
Based on Theorem 5, we note that $2\eta C$ is a small constant. When $T$ is sufficiently large such that the model $\mathbf{w}_{E,B}^{-U}$ is the empirical risk minimizer $\hat{\mathbf{w}}^{-U}$ of the objective function $F_{\mathcal{S}\backslash U}(\mathbf{w})$, the last two terms thus approach 0. Therefore, we have:
$F(\tilde{\mathbf{w}}_{E,B}^{-U})-\mathbb{E}[F\left(\mathbf{w}^{*}\right)]\leq\mathcal{O}\left(\frac{4L^{2}}{\lambda(n-m)}\right).$ (66)
where $\tilde{\mathbf{w}}_{E,B}^{-U}$ denotes the output point $\bar{\Omega}(\Omega(\mathcal{S});T(\mathcal{S}))$ and $\mathbf{w}^{*}$ is the minimizer of the population risk in (1). Therefore, the above upper bound on the excess risk implies that we can delete at least
$m=n-\frac{4L^{2}}{\gamma\lambda}$ (67)
samples while still ensuring an excess risk guarantee of $\gamma$. Intuitively, according to Theorem 5, in the optimal situation, we discuss the excess loss of our algorithm and demonstrated its superiority over previous methods, indicating improved deletion capacity. In particular, Sekhari et al. [2] derive the following generalization bound, where $\mathbf{w}_{E,B}$ is required to be the empirical risk minimizer:
$F(\tilde{\mathbf{w}}_{E,B}^{-U})-\mathbb{E}[F\left(\mathbf{w}^{*}\right)]\leq\mathcal{O}\left(\frac{mL^{2}}{\lambda(n-m)}+\sqrt{d}\sigma L+\frac{4L^{2}}{\lambda n}\right),$ (68)
where $\sigma$ represents a term associated with noise, while $d$ denotes the dimensionality of the model.
Using the fact that $n=\Omega(m)$, it is noted that in (66), our generalization bound is a constant term of $\frac{L^{2}}{\lambda(n-m)}$, which was omitted in previous works. This implies that, compared to previous works, our method achieves better generalization results and improved deletion capacity under same assumption (e.g. $\lambda$-strongly convex, $L$-Lipschitz, and $M$-Smoothness of the objective function, as well as the model being the empirical risk minimizer). ∎
E.2 Limitation and Future Work
Compared to prior certified unlearning works, our paper is the first to achieve near-instantaneous certified unlearning under non-convex non-smooth settings (Theorem 1), and under the same convex and smooth assumptions, our method provides better generalization, deletion capacity, and unlearning guarantees (Theorem 5) than previous works. Moreover, we theoretically (Subsection 4.4) and experimentally (Subsection C.2) demonstrate that our method is more efficient in terms of precomputation/storage and unlearning time complexity than previous approaches.
Although we have made the above contributions, there are still some significant limitations:
  1. 1.
    Firstly, our proposed algorithm necessitates $\mathcal{O}(nTd)$ and $\mathcal{O}(nd)$, which remains costly, particularly in scenarios involving extensive data and larger-scale models. Consequently, conducting experiments concurrently with both large-scale datasets and models is anticipated to significantly extend the time required for precomputing unlearning statistics.
  2. 2.
    In addition, although we establish the analysis of certified unlearning for non-convex scenarios, as demonstrated by the results in Theorem 4, it is important to recognize that the proposed method is highly sensitive to both the experiment’s hyperparameters and the stability of the training process in non-convex settings. If the training process is unstable, our method will obtain a significant approximation error. For example, in Subsection D.1 (in Table 9), when using an aggressive step size (e.g. step size increasing from 0.1 to 0.5), the approximation error becomes very large (with results to the power of 7).
Future work. Therefore, in order to solve the problems mentioned above, we briefly propose some solutions. Specifically, our proposed framework offers the feasibility to unlearn an arbitrary subset of the entire datasets, allowing a trade-off that further reduces the storage complexity, for instance, consider a typical federated unlearning scenario at the user level, where there are $k$ users ($k<<n$). By focusing on the unlearning of a single user’s dataset or a subset of users’ datasets, we can maintain $k$ vectors for $k$ users’ dataset instead of maintaining $n$ vectors for $n$ all data points. Therefore, we can further reduce storage to $\mathcal{O}(kd)$. Besides, batching is also beneficial for GPU parallel computation.
Furthermore, based on our analysis in Theorem 4 and Subsection D.1, the main source of the approximation error is $\rho$, which mainly consists of two components: (1) the step size $\eta_{e,b}$ and (2) the maximum eigenvalue $\lambda_{1}$ of Hessian $\frac{1}{|\mathcal{B}_{e,b}|}\mathbf{H}_{e,b}$. Therefore, effectively reducing $\lambda_{1}$ during optimization, so that the training enters a flatter region, will not only improve the generalization ability of learned model but also reduce the approximation error in the unlearning stage. A possible approach is through sharpness-aware minimization [55, 56], which can effectively reduce $\lambda_{1}$.
E.3 Online Unlearning-Repairing Strategy for Algorithm 1
While our approach does not require accessing any datasets in Stage III, we can mitigate the performance degradation caused by excessive deletions and make our algorithm more robust through a simple fine-tuning strategy if we can still access the remaining dataset in Stage III.
Specifically, when the continuous incoming deletion requests to forget data, we can rapidly implement certified unlearning through vector addition. This involves augmenting the current model with approximators, where errors accumulate during this process and thus cause performance degradation. Suppose, upon the arrival of a new round of deletion requests for a particular model and dataset, the model’s accuracy significantly drops due to the removal of a substantial amount of data. In such the scenario, we can address this performance degradation by fine-tuning the current model for $E_{r}$ epochs on the remaining dataset, referred to as the ‘repairing’ process. Additionally, by computing the approximators for the remaining dataset during the repairing process, and adding these approximators during the repairing process to the pre-stored approximators during the learning process, we obtain the new set of approximators for the remaining data.
1 Stage I: Learning model $\mathbf{w}_{E,B}$ on dataset $S=\{z_{i}\}_{i=1}^{n}$:
2 for  $e=0,1\cdots,E$ do
3       for $b=0,1\cdots,B$ do
4             Compute gradient: $\mathbf{g}\!\leftarrow\!\frac{1}{|\mathcal{B}_{e,b}|}\sum_{i\in\mathcal{B}_{e,b}}\!\!\!\nabla\ell\left(\mathbf{w}_{e,b};z_{i}\right)\times\max(1,\frac{C}{\left\|{\nabla\ell\left(\mathbf{w}_{e,b};z_{i}\right)}\right\|})$,
6       end for
7
8 end for
9
10Stage II: Pre-computing and Pre-storing statistics $\mathcal{T}(\mathcal{S})=\{\mathbf{a}_{E,B}^{-u_{j}}\}_{j=1}^{n}$:
11 for  $j=1,2\cdots,n$ do
12       Recursive computation by using HVP: $\mathbf{a}_{E,B}^{-u_{j}}\leftarrow\sum_{e=1}^{E}\mathbf{M}_{e,b(u_{j})}\nabla\ell(\mathbf{w}_{e,b(u_{j})};u_{j})$,
13       where $\mathbf{M}_{e,b(u_{j})}=\frac{\eta_{e,b(u_{j})}}{|\mathcal{B}_{e,b(u_{j})}|}\prod^{E}_{k=e}\prod_{b=b(u_{j})+1}^{B-1}\!(\mathbf{I}-\frac{\eta_{k,b}}{|\mathcal{B}_{k,b}|}\mathbf{H}_{k,b})$
14 end for
15
16Stage III: Unlearning when user requests to forget the subset $U$=$\{u_{j}\}_{j=1}^{m}$ on model $\mathbf{w}_{E,B}$:
17 Compute: $\bar{\mathbf{w}}_{E,B}^{-U}\!\leftarrow\!\mathbf{w}_{E,B}\!+\!\mathbf{a}_{E,B}^{-U},$ $c\!\leftarrow\!{2\eta C}\frac{\sqrt{2\!\ln\!{(1.25/\delta})}}{\epsilon}{\zeta_{T}^{-U}}\!\!$    Delete statistics: $\{\mathbf{a}_{E,B}^{-u_{j}}\}_{j=1}^{m}$
18 Sample: $\sigma\sim\mathcal{N}(0,c\mathbf{I})$, Return: $\tilde{\mathbf{w}}_{E,B}^{-U}=\bar{\mathbf{w}}_{E,B}^{-U}+\sigma$
19Reparing (Fine-tuning) the model $\tilde{\mathbf{w}}^{-U}_{E,B}$ on the remaining dataset $S\backslash U=\{z_{r}\}_{r=1}^{n-m}$:
20 for $e=0,...,E_{r}$ do
21       for $b=0,1\cdots,B$ do
22             Compute gradient: $\mathbf{g}\!\leftarrow\!\frac{1}{|\mathcal{B}_{e,b}|}\sum_{r\in\mathcal{B}_{e,b}}\!\!\!\nabla\ell\left(\mathbf{w}_{e,b};z_{r}\right)\times\max(1,\frac{C}{\left\|{\nabla\ell\left(\mathbf{w}_{e,b};z_{r}\right)}\right\|})$,
24       end for
25
26 end for
27
28Pre-computing statistics $\mathcal{T}(\mathcal{S})=\{\mathbf{a}_{E,B}^{-u_{r}}\}_{r=1}^{n-m}$ of the remaining dataset $S\backslash U=\{z_{r}\}_{r=1}^{n-m}$:
29 for  $r=1,2\cdots,{n-m}$ do
30       Recursive computation by using HVP:
31       $\mathbf{a}_{E+E_{r},B}^{-u_{r}}\leftarrow\mathbf{M}_{e,b(u_{r})}\mathbf{a}_{E,B}^{-u_{r}}+\sum_{e=1}^{E_{r}}\mathbf{M}_{e,b(u_{r})}\nabla\ell(\mathbf{w}_{e,b(u_{r})};u_{r})$
32       where: $\mathbf{M}_{e,b(u_{r})}=\frac{\eta_{e,b(u_{r})}}{|\mathcal{B}_{e,b(u_{r})}|}\prod^{E_{r}}_{k=e}\prod_{b=b(u_{r})+1}^{B-1}(\mathbf{I}-\frac{\eta_{k,b}}{|\mathcal{B}_{k,b}|}\mathbf{H}_{k,b})$
33 end for
34
Algorithm 2 Hessian-free Online Unlearning-Repairing Strategy
E.4 Social Impacts
This paper introduces a Hessian-free online unlearning method via maintaining a statistical vector for each data sample, computed through affine stochastic recursion approximation of the difference in the retrained and the unlearn models. We have shown that our approach outperforms the existing results in terms of time/memory costs, while also enhancing generalization. While the unlearning algorithm was initially proposed to assist organizations in complying with data protection regulations (e.g., GDPR) and help build trust in AI systems by demonstrating the ability to correct mistakes or biases, it is important to be mindful of potential negative consequences: Unlearning algorithms can introduce unforeseen biases or inaccuracies in subsequent predictions or decisions. In addition, adversarial actors could potentially exploit machine unlearning processes to manipulate or subvert models.

### v4

@@ E.1 Impact of Step Sizes
Figure 10: Impact of Epoch. (a) (b) illustrate that in the distance for LR and CNN by varying epoch from 1 and 100. As shown in the Figure (a) (b), the error accumulates with the increase in epochs. However, with an appropriate choice of step size, such growth is acceptable (Seed 42).
E.2 Impact of Epochs
In this subsection, we began investigating the distance metric as the number of epochs increases.
Configurations: We investigated the distance by varying the epoch from 1 to 100. We conducted experiments with a high deletion rate of 20% and a low deletion rate of 5%. The selection of other hyperparameters remains consistent with the ones used previously.
Impact of epochs. As shown in Figure 10, the distance exhibits convergent behavior, ultimately stabilizing at a constant value. This phenomenon is consistent with our theoretical findings, as stated in Theorem 4, which demonstrates that the approximation error decreases with an increasing number of epochs for convex models. The contractive mapping ensures a gradual reduction in the final error. In the case of non-convex models, the error tends to increase with the number of epochs, ultimately stabilizing, a phenomenon that can be explained by the degeneracy of the Hessian in over-parameterized models (Sagun et al. (2018)), as we will discuss in further detail later.
E.3 Impact of Stepsize Decay
These experiments aim to study the impacts of stepsize decay. We consider two scenarios: one with stepsize decay ($q=0.995$) and the other without stepsize decay ($q=1$).
Impact of decay rates of stepsize: We can observe from Figure 11 that the Euclidean distance for the stepsize decay strategy is generally smaller compared to that without the stepsize decay. This aligns with our earlier analysis of step size in Subsection E.1, i.e., a smaller step size leads to a smaller approximation error, and choosing an appropriate stepsize decay rate ensures a continuous reduction in step size, preventing the unbounded growth of the approximation error.
Additionally, as in Figure 11 (b), It can be observed that the distance metric tends to stabilize even without stepsize decay. This result can be explained by the conclusions in Sagun et al. (2018; 2016). Specifically, Sagun et al. (2018; 2016) analyzed the spectral distribution of the Hessian eigenvalues during CNN training. It reveals that, initially, the Hessian eigenvalues are symmetrically distributed around 0. As training progresses, they converge towards 0, indicating Hessian degeneracy. Toward the end of training, only a small number of eigenvalues become negative, and a large portion of the Hessian eigenvalues approach 0 when the training approaches convergence. Recall that our previous analysis of the stepsize in Subsection E.1, where we demonstrate that the approximation error is mainly determined by $\rho$ of $\mathbf{I}-{\eta_{e,b}}\mathbf{H}_{e,b}$, and $\rho$ is determined by $\eta_{e,b}$ and $\lambda_{1}$. As the number of iterations increases, the eigenvalues of the Hessian $\mathbf{H}_{e,b}$ tends to 0. This implies a decrease in $\lambda_{1}$, which leads to a slower increase in error. Therefore, in Figure 11 (b), the distance gradually stabilizes even without a decay strategy due to the decrease in $\lambda_{1}$ during training, and it still increases due to the existence of some larger outlier eigenvalues.
@@ E.3 Impact of Stepsize Decay
Figure 11: Impacts of the stepsize decay rate on the Euclidean distances for (a) the LR and (b) the CNN with 20% data to be forgotten during 25 epochs. The red line represents learning with a decay rate of 0.995, while the blue line represents without decay. It can be observed that stepsize decay effectively reduces the approximation error of the unlearned model. (Seed 42)
E.4 Impact of Gradient Clipping
Stochastic gradient algorithms are often unstable when applied to functions that lack Lipschitz continuity and/or bounded gradients. It is well-established that exploding gradients pose a significant challenge in deep learning applications (Hanin (2018)). Since our proposed method relies on backtracking historical gradients, the approach may fail completely when gradients tend toward infinity. To circumvent this issue, we incorporated gradient clipping into all the aforementioned experiments. Gradient clipping is a widely adopted technique that ensures favorable performance in deep network training by restricting gradient norms from becoming excessively large (Mai & Johansson (2021); Chen et al. (2020); Zhang et al. (2020)). However, this operation indicates that the gradients in the Taylor expansion of Eq. 2 are not the true gradients but rather the clipped gradients, which introduces an approximation error. Our subsequent experiments will demonstrate that this error can be rendered negligible. It is worth noting that our method does not inherently rely on gradient clipping but instead imposes requirements on the stability of the learning process.
Configurations: We investigate the impact of gradient clipping with thresholds set at {0.3, 0.5, 0.8, 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 15, 20, 25, 30, 35}. We remove 20% of the data while keeping other hyperparameter settings consistent with those described in the main text.
Figure 12: Impacts of the gradient clipping on the distance metric for the LR (orange) and the CNN (blue) with 20% data to be forgotten. (Seed 42)
Impact of gradient clipping. As evidenced in Figure 12, the implementation of small clipping thresholds results in a scaling down of the gradient proportional to its norm, producing an effect similar to using a smaller step size. Consequently, the application of lower clipping thresholds maintains the error at a relatively minimal level. However, as the gradient clipping threshold surpasses a certain range, the error introduced by the clipping process becomes the predominant factor, culminating in an escalation of the overall error. Nonetheless, most gradients no longer require clipping during the optimization process. Therefore, the induced error diminishes until the threshold reaches a sufficiently elevated value, at which point no supplementary error is introduced. Once this threshold is attained, gradient clipping ceases to exert an influence on the performance of our proposed unlearning algorithm.